Show step by step:

Find the inverse for y= 2x2-8

1
Expert's answer
2021-03-25T08:41:44-0400

"y = 2x^2 - 8"


"y + 8 = 2x^2"


"x^2 = {y+8 \above{1pt} 2}"


"x = \sqrt{\smash[b]{y+8 \above{1pt} 2}}"